I forget how to factor. Can you please help:
3(x+4)^2=8
1
Expert's answer
2012-09-04T07:10:21-0400
3(x+4)^2=8
3(x^2+8x+16)=8
3x^2+24x+48=8
3x^2+24x+40=0
D/4=144-40*3=24
x1,2=(-12+-sqrt(24))/3
x1=(-12+sqrt(24))/3
x2=(-12-sqrt(24))/3
Therefore due to factor theorem 3x^2+24x+40=3(x-x1)(x-x2)